Grade 12Mechanics

A block of mass 1kg is pushed against a rough vertical wall with a force of 20N , coefficient of static friction being 1/4 . Another horizontal force of 10N is applied on the block in a direction parallel to the wall. With what acceleration will the block move?

Shanu
8 Years agoGrade 12

1 Answer

Khimraj
8 Years ago
net force on block due to gravity and horizontal force = \sqrt{(mg)^{2} + 15^{2}} = 5\sqrt{13}N
friction force = ¼*20 = 5
so acceleration = (5\sqrt{13} – 5)/1 = 5(\sqrt{13} – 1) m/s2
Hope it clears